### v1

@@ 3 Ensemble Transport Filter
In this section, a transport map is constructed to approximate the transition probability for the analysis step. In Section 3.1, we formulate the construction as an optimization task on a Maximum Mean Discrepancy loss by matching the posterior expectation. The algorithm of the ensemble-based transport filter is presented in Section 3.2. Finally, in Section 3.3, a variance penalty is introduced to improve the robustness of MMD. Moreover, an linear transport filter is derived, and it shares the similar form to EnKF.
@@ 3.1 Proposal transition generated by transport map in analysis step
In filtering, we typically use the posterior mean as the state estimate. Here, we start by considering the posterior mean of an arbitrary function $f\in\mathcal{C}(\mathbb{R}^{n})$. For simplicity, we suppress the subscripts of the probability density functions in the derivation.
(7) $\mathbb{E}_{\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}}\left[f(\boldsymbol{x}_{k})\right]$ $=\int f(\boldsymbol{x}_{k})\pi(\boldsymbol{x}_{k}\mid\boldsymbol{y}_{1:k})\mathrm{d}\boldsymbol{x}_{k}$
$=\frac{1}{A}\int f(\boldsymbol{x}_{k})\pi(\boldsymbol{y}_{k}\mid\boldsymbol{x}_{k})\pi(\boldsymbol{x}_{k}\mid\boldsymbol{y}_{1:k-1})\mathrm{d}\boldsymbol{x}_{k}$
$=\frac{1}{A}\int f(\boldsymbol{x}_{k})\pi(\boldsymbol{y}_{k}\mid\boldsymbol{x}_{k})\pi(\boldsymbol{x}_{k}\mid\boldsymbol{x}_{k-1})\pi(\boldsymbol{x}_{k-1}\mid\boldsymbol{y}_{1:k-1})\mathrm{d}\boldsymbol{x}_{k}\mathrm{d}\boldsymbol{x}_{k-1},$
@@
in which Bayes’ theorem and Chapman-Kolomogrov equation are used, and $A$ is a normalization factor.
In the particle filter, the posterior distribution at the previous time step can be approximated by the summation of Dirac-delta centered at particles with equal weights, i.e.,
(8) $\pi(\boldsymbol{x}_{k-1}\mid\boldsymbol{y}_{1:k-1})\approx\frac{1}{N}\sum_{i=1}^{N}\delta_{\boldsymbol{x}_{k-1}^{(i)}}(\boldsymbol{x}_{k-1}).$
Then, the samples of forecast distribution $\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k-1}}$ can be draw from the transition density $\pi(\boldsymbol{x}_{k}\mid\boldsymbol{x}_{k-1})$ via the discrete dynamical model Eq. 1, which gives $\{\hat{\boldsymbol{x}}_{k}^{(i)}\}_{i=1}^{N}$. Filtering prior distribution at time step $k$ can be estimated by
(9) $\pi(\boldsymbol{x}_{k}\mid\boldsymbol{y}_{1:k-1})\approx\frac{1}{N}\sum_{i=1}^{N}\delta_{\hat{\boldsymbol{x}}_{k}^{(i)}}(\boldsymbol{x}_{k-1}).$
Observations are introduced through particle weights computed with the likelihood function. The posterior expectation of function $f(\boldsymbol{x})$ can be estimated by
(10) $\mathbb{E}_{\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}}\left[f(\boldsymbol{x}_{k})\right]\approx\sum_{i=1}^{N}w_{i}f(\hat{\boldsymbol{x}}_{k}^{i}),\quad w_{i}=\frac{\pi(\boldsymbol{y}_{k}\mid\hat{\boldsymbol{x}}_{k}^{(i)})}{\sum_{j=1}^{N}\pi(\boldsymbol{y}_{k}\mid\hat{\boldsymbol{x}}_{k}^{(j)})}.$
Combining Eq. 5-Eq. 6, the posterior expectation Eq. 7 can also be represented by
$\mathbb{E}_{\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}}\left[f(\boldsymbol{x}_{k})\right]=\int f(\boldsymbol{x}_{k})\pi(\boldsymbol{x}_{k}\mid\hat{\boldsymbol{x}}_{k},\boldsymbol{y}_{k})\pi(\hat{\boldsymbol{x}}_{k}\mid\boldsymbol{x}_{k-1})\pi(\boldsymbol{x}_{k-1}\mid\boldsymbol{y}_{1:k-1})\mathrm{d}\hat{\boldsymbol{x}}_{k}\mathrm{d}\boldsymbol{x}_{k-1}\mathrm{d}\boldsymbol{x}_{k}.$
@@
As discussed in Section 2.2, the transition probability density $\pi_{\boldsymbol{X}_{k}\mid\hat{\boldsymbol{X}}_{k},\boldsymbol{Y}_{k}}$ is implicitly defined through the likelihood function. Here, we find a transition probability density $q^{\mathcal{T}_{k}}_{\boldsymbol{X}_{k}\mid\hat{\boldsymbol{X}}_{k},\boldsymbol{y}_{k}}$ through a parametric transport map $\mathcal{T}_{k}$ to obtain particles from the posterior distribution directly. To eliminate the intermediate variable $\hat{\boldsymbol{X}}_{k}$, we denote
$q^{\mathcal{T}_{k}}(\boldsymbol{x}_{k}\mid\boldsymbol{x}_{k-1},\boldsymbol{y}_{k})=\int q^{\mathcal{T}_{k}}(\boldsymbol{x}_{k}\mid\hat{\boldsymbol{x}}_{k},\boldsymbol{y}_{k})\pi(\hat{\boldsymbol{x}}_{k}\mid\boldsymbol{x}_{k-1})\mathrm{d}\hat{\boldsymbol{x}}_{k}.$
Insert $q^{\mathcal{T}_{k}}(\boldsymbol{x}_{k}\mid\boldsymbol{x}_{k-1},\boldsymbol{y}_{k})$ into Eq. 7, we can rewrite it as
(11) $\mathbb{E}_{\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}}\left[f(\boldsymbol{x}_{k})\right]$ $=$
$\int f(\boldsymbol{x}_{k})$ $h^{\mathcal{T}_{k}}(\boldsymbol{x}_{k-1:k},\boldsymbol{y}_{k})q^{\mathcal{T}_{k}}(\boldsymbol{x}_{k}\mid\boldsymbol{x}_{k-1},\boldsymbol{y}_{k})\pi(\boldsymbol{x}_{k-1}\mid\boldsymbol{y}_{1:k-1})\mathrm{d}\boldsymbol{x}_{k}\mathrm{d}\boldsymbol{x}_{k-1},$
where $h^{\mathcal{T}_{k}}(\boldsymbol{x}_{k-1:k},\boldsymbol{y}_{k})=\frac{\pi(\boldsymbol{y}_{k}\mid\boldsymbol{x}_{k})\pi(\boldsymbol{x}_{k}\mid\boldsymbol{x}_{k-1})}{Aq^{\mathcal{T}_{k}}(\boldsymbol{x}_{k}\mid\boldsymbol{x}_{k-1},\boldsymbol{y}_{k})}$.
@@
If a transport map $\mathcal{T}_{k}$ is selected, the transition probability density $q^{\mathcal{T}_{k}}_{\boldsymbol{X}_{k}\mid\hat{\boldsymbol{X}}_{k},\boldsymbol{y}_{k}}$ corresponding to the map satisfy
(12) $h^{\mathcal{T}_{k}}(\boldsymbol{x}_{k-1:k},\boldsymbol{y}_{k})\equiv 1.$
Thus, we obtain a transport map that transports the prior samples to the posterior samples. The posterior expectation of function $f(\boldsymbol{x})$ can be estimated by
$\mathbb{E}_{\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}}\left[f(\boldsymbol{x}_{k})\right]\approx\frac{1}{N}\sum_{i=1}^{N}f(\mathcal{T}_{k}(\boldsymbol{x}_{k}^{(i)})).$
@@
It is difficult to directly solving Eq. 12 because the normalization constant $A$ is inaccessible. We can not obtain an analytical form for the transition probability corresponding to the nonlinear transport map. However, we can compute $\mathcal{T}_{k}$ indirectly. In a compact form, we can use the Chapman-Kolmogorov formula to obtain an approximated posterior probability density,
$q^{\mathcal{T}_{k}}(\boldsymbol{x}_{k}\mid\boldsymbol{y}_{1:k})=\int q^{\mathcal{T}_{k}}(\boldsymbol{x}_{k}\mid\boldsymbol{x}_{k-1},\boldsymbol{y}_{k})\pi(\boldsymbol{x}_{k-1}\mid\boldsymbol{y}_{1:k-1})\mathrm{d}\boldsymbol{x}_{k-1}.$
If the approximated posterior $q^{\mathcal{T}_{k}}(\boldsymbol{x}_{k}\mid\boldsymbol{y}_{1:k})$ is equal to the true posterior, the desired transport map $\mathcal{T}_{k}$ is found.
@@
(Lemma 9.3.2 in [26]) Let $(\Omega,d)$ be a metric space, and $p,q$ be two probability measure defined on $\Omega$. Then $p=q$ if and only if $\mathbb{E}_{\boldsymbol{x}\sim p}[f(\boldsymbol{x})]=\mathbb{E}_{\boldsymbol{y}\sim q}[f(\boldsymbol{y})]$ for all $f\in\mathcal{C}(\Omega)$, where $\mathcal{C}(\Omega)$ is the space of bounded continuous on $\Omega$.
The following proposition is a direct result of Lemma 3.1, and it proves that the resulting posterior probability density gives an approximation of the true posterior probability density under certain conditions.
@@
Given transport map $\mathcal{T}_{k}:\mathbb{R}^{n}\to\mathbb{R}^{n}$, the transition probability density $q^{\mathcal{T}_{k}}_{\boldsymbol{X}_{k}\mid\hat{\boldsymbol{X}}_{k},\boldsymbol{y}_{k}}$ corresponds to transport map $\mathcal{T}_{k}$ generates the posterior $q^{\mathcal{T}_{k}}_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}$. If the expectation of posterior and true posterior are matched for all $\ f\in\mathcal{C}(\mathbb{R}^{n})$,
$\int f(\boldsymbol{x}_{k})\pi(\boldsymbol{x}_{k}\mid\boldsymbol{y}_{1:k})\mathrm{d}\boldsymbol{x}_{k}=\int f(\boldsymbol{x}_{k})q^{\mathcal{T}_{k}}(\boldsymbol{x}_{k}\mid\boldsymbol{y}_{1:k})\mathrm{d}\boldsymbol{x}_{k},$
then the approximated posterior $q^{\mathcal{T}_{k}}(\boldsymbol{x}_{k}\mid\boldsymbol{y}_{1:k})$ is identical to the true posterior.
Based on Proposition 3.2, we can reformulate the problem of finding transport map $\mathcal{T}_{k}$ as the optimization probem,
(13) $\mathcal{T}_{k}=\underset{\mathcal{S}}{\arg\min}\ \underset{f\in\mathcal{C}(\mathbb{R}^{n})}{\sup}\left(\mathbb{E}_{\boldsymbol{x}\sim\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}}[f(\boldsymbol{x})]-\mathbb{E}_{\boldsymbol{y}\sim\pi^{\mathcal{S}}_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}}[f(\boldsymbol{y})]\right).$
It is challenging to solve the optimization on the continuous function space $\mathcal{C}(\Omega)$. To make it efficient and feasible, we select an RKHS space $\mathcal{H}$ defined on a compact metric space $\Omega$ with reproducing kernel $k(\cdot,\cdot)$. When the kernel of $\mathcal{H}$ is universal, the properties of the Lemma 3.1 are held. Let $\mathcal{F}$ be a unit ball in RKHS $\mathcal{H}$. Then $\operatorname{MMD}[p,q;\mathcal{F}]=0$ if and only if $p=q$. The target of the optimization problem Eq. 13 can be rewritten as
(14) $\underset{f\in\mathcal{F}}{\sup}\left(\mathbb{E}_{\boldsymbol{x}\sim\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}}[f(\boldsymbol{x})]-\mathbb{E}_{\boldsymbol{y}\sim\pi^{\mathcal{S}}_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}}[f(\boldsymbol{y})]\right).\$
The formula Eq. 14 is also called Maximum Mean Discrepancy(MMD). The MMD is a reasonable measure of the difference between two distributions. When MMD equals zero, two distributions are identical. With the reproducing property of RKHS, we do not need to compute the expectation of a specific function $f$ directly. It suffices to compute Eq. 14 by Eq. 15 in the following lemma.
@@
(Lemma 5 in [11]) Given $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$ independent variables with distribution $p$ , $\boldsymbol{y}$ and $\boldsymbol{y}^{\prime}$ independent variables with distribution $q$. Assuming that the expectation of feature mapping $\phi(\cdot)$ of RKHS $\mathcal{H}$ exists. Then
(15) $\operatorname{MMD}\limits^{2}[p,q;\mathcal{F}]=\mathbb{E}_{\boldsymbol{x},\boldsymbol{x}^{\prime}\sim p}[k(\boldsymbol{x},\boldsymbol{x}^{\prime})]-2\mathbb{E}_{\boldsymbol{x}\sim p,\boldsymbol{y}\sim q}[k(\boldsymbol{x},\boldsymbol{y})]+\mathbb{E}_{\boldsymbol{y},\boldsymbol{y}^{\prime}\sim p}[k(\boldsymbol{y},\boldsymbol{y}^{\prime})].$

### v2

@@ 3 Ensemble Transport Filter
In this section, we construct a transport map to approximate the transition probability for the analysis step. Section 3.1 formulates this construction as an optimization problem based on the Maximum Mean Discrepancy loss by matching the posterior expectation. The algorithm for the ensemble transport filter is detailed in Section 3.2. To enhance the robustness of MMD, a variance penalty is introduced, and then an extension for high-dimensional problems is discussed. Finally, a linear version of the transport filter is derived, which takes a form similar to that of the Ensemble Kalman Filter.
@@ 3.1 Proposal transition generated by transport map in analysis step
In filtering, we typically use the posterior mean as the state estimate. Here, we start by considering the posterior expectation of an arbitrary function $f\in\mathcal{C}(\mathbb{R}^{n})$. For simplicity, we suppress the subscripts of the probability density functions in the derivation.
$\mathbb{E}_{\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k}}}\left[f(\bm{x}_{k})\right]$ $=\int f(\bm{x}_{k})\pi(\bm{x}_{k}\mid\bm{y}_{1:k})\mathrm{d}\bm{x}_{k}$ (3.7)
$=\frac{1}{A}\int f(\bm{x}_{k})\pi(\bm{y}_{k}\mid\bm{x}_{k})\pi(\bm{x}_{k}\mid\bm{y}_{1:k-1})\mathrm{d}\bm{x}_{k}$
$=\frac{1}{A}\int f(\bm{x}_{k})\pi(\bm{y}_{k}\mid\bm{x}_{k})\pi(\bm{x}_{k}\mid\bm{x}_{k-1})\pi(\bm{x}_{k-1}\mid\bm{y}_{1:k-1})\mathrm{d}\bm{x}_{k}\mathrm{d}\bm{x}_{k-1},$
@@
in which Bayes’ theorem and Chapman-Kolomogrov equation are used, and $A=\int\pi(\bm{y}_{k}\mid\bm{x}_{k})\pi(\bm{x}_{k}\mid\bm{y}_{1:k-1})\mathrm{d}\bm{x}_{k}$ is a normalization factor.
Combining (2.5)-(2.6), the posterior expectation (3.7) can also be represented by
$\mathbb{E}_{\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k}}}\left[f(\bm{x}_{k})\right]=\int f(\bm{x}_{k})\pi(\bm{x}_{k}\mid\hat{\bm{x}}_{k},\bm{y}_{k})\pi(\hat{\bm{x}}_{k}\mid\bm{x}_{k-1})\pi(\bm{x}_{k-1}\mid\bm{y}_{1:k-1})\mathrm{d}\hat{\bm{x}}_{k}\mathrm{d}\bm{x}_{k-1}\mathrm{d}\bm{x}_{k}.$
@@
As discussed in 2.2, the transition probability density $\pi_{\bm{X}_{k}\mid\hat{\bm{X}}_{k},\bm{Y}_{k}}$ is implicitly defined through the likelihood function. Here, we aim to find a transition probability density $q^{\mathcal{T}_{k}}_{\bm{X}_{k}\mid\hat{\bm{X}}_{k},\bm{y}_{k}}$ through a parametric transport map $\mathcal{T}_{k}$ to obtain particles from the posterior distribution directly. To eliminate the intermediate variable $\hat{\bm{X}}_{k}$, we denote
$q^{\mathcal{T}_{k}}(\bm{x}_{k}\mid\bm{x}_{k-1},\bm{y}_{k})=\int q^{\mathcal{T}_{k}}(\bm{x}_{k}\mid\hat{\bm{x}}_{k},\bm{y}_{k})\pi(\hat{\bm{x}}_{k}\mid\bm{x}_{k-1})\mathrm{d}\hat{\bm{x}}_{k}.$
Then we have
$\mathbb{E}_{\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k}}}\left[f(\bm{x}_{k})\right]=\int f(\bm{x}_{k})q^{\mathcal{T}_{k}}(\bm{x}_{k}\mid\bm{x}_{k-1},\bm{y}_{k})\pi(\bm{x}_{k-1}\mid\bm{y}_{1:k-1})\mathrm{d}\bm{x}_{k}\mathrm{d}\bm{x}_{k-1}.$
Moreover, insert $q^{\mathcal{T}_{k}}(\bm{x}_{k}\mid\bm{x}_{k-1},\bm{y}_{k})$ into (3.7), we can rewrite it as
$\mathbb{E}_{\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k}}}\left[f(\bm{x}_{k})\right]$ $=$ (3.8)
$\int f(\bm{x}_{k})$ $h^{\mathcal{T}_{k}}(\bm{x}_{k-1:k},\bm{y}_{k})q^{\mathcal{T}_{k}}(\bm{x}_{k}\mid\bm{x}_{k-1},\bm{y}_{k})\pi(\bm{x}_{k-1}\mid\bm{y}_{1:k-1})\mathrm{d}\bm{x}_{k}\mathrm{d}\bm{x}_{k-1},$
where $h^{\mathcal{T}_{k}}(\bm{x}_{k-1:k},\bm{y}_{k})=\frac{\pi(\bm{y}_{k}\mid\bm{x}_{k})\pi(\bm{x}_{k}\mid\bm{x}_{k-1})}{Aq^{\mathcal{T}_{k}}(\bm{x}_{k}\mid\bm{x}_{k-1},\bm{y}_{k})}$.
@@
If a suitable transport map $\mathcal{T}_{k}$ has been chosen, the transition probability density $q^{\mathcal{T}_{k}}_{\bm{X}_{k}\mid\hat{\bm{X}}_{k},\bm{y}_{k}}$ corresponding to the transport map should satisfy
$h^{\mathcal{T}_{k}}(\bm{x}_{k-1:k},\bm{y}_{k})\equiv 1.$ (3.9)
@@
Directly solving (3.9) is challenging, as the normalization constant $A$ is generally inaccessible. Furthermore, an analytical form for the transition probability corresponding to the nonlinear transport map cannot be derived. However, the transport map $\mathcal{T}_{k}$ can be computed indirectly.
In a compact form, we can use the Chapman-Kolmogorov formula to obtain an transported posterior probability density,
$q^{\mathcal{T}_{k}}(\bm{x}_{k}\mid\bm{y}_{1:k})=\int q^{\mathcal{T}_{k}}(\bm{x}_{k}\mid\bm{x}_{k-1},\bm{y}_{k})\pi(\bm{x}_{k-1}\mid\bm{y}_{1:k-1})\mathrm{d}\bm{x}_{k-1}.$
Next, we construct the desired transport map indirectly by matching posterior expectations. To begin, we introduce the following lemma.
@@
(Lemma 9.3.2 in [26]) Let $(\Omega,d)$ be a metric space, and $p,q$ be two probability measure defined on $\Omega$. Then $p=q$ if and only if $\mathbb{E}_{\bm{x}\sim p}[f(\bm{x})]=\mathbb{E}_{\bm{y}\sim q}[f(\bm{y})]$ for all $f\in\mathcal{C}(\Omega)$, where $\mathcal{C}(\Omega)$ is the space of bounded continuous on $\Omega$.
The following proposition is a direct result of Lemma 3.1, and it suggests that the transported posterior $q^{\mathcal{T}_{k}}(\bm{x}_{k}\mid\bm{y}_{1:k})$ gives an approximation of the true posterior probability density under certain conditions.
@@
Given transport map $\mathcal{T}_{k}:\mathbb{R}^{n}\to\mathbb{R}^{n}$, if the expectation of posterior and true posterior are matched for all $f\in\mathcal{C}(\mathbb{R}^{n})$,
$\int f(\bm{x}_{k})\pi(\bm{x}_{k}\mid\bm{y}_{1:k})\mathrm{d}\bm{x}_{k}=\int f(\bm{x}_{k})q^{\mathcal{T}_{k}}(\bm{x}_{k}\mid\bm{y}_{1:k})\mathrm{d}\bm{x}_{k},$
then the approximated posterior $q^{\mathcal{T}_{k}}(\bm{x}_{k}\mid\bm{y}_{1:k})$ is identical to the true posterior and $\mathcal{T}_{k}$ is the desired transport map that transfer prior to posterior.
Based on Propsition 3.2, we can reformulate the problem of finding transport map $\mathcal{T}_{k}$ as the optimization probem,
$\mathcal{T}_{k}=\underset{\mathcal{S}}{\arg\min}\ \underset{f\in\mathcal{C}(\mathbb{R}^{n})}{\sup}\left(\mathbb{E}_{\bm{x}\sim\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k}}}[f(\bm{x})]-\mathbb{E}_{\bm{y}\sim\pi^{\mathcal{S}}_{\bm{X}_{k}\mid\bm{Y}_{1:k}}}[f(\bm{y})]\right),$ (3.10)
where $\pi^{\mathcal{S}}_{\bm{X}_{k}\mid\bm{Y}_{1:k}}$ is the approximated posterior given by transport map $\mathcal{S}$. It is challenging to solve the optimization on the continuous function space $\mathcal{C}(\Omega)$. To make it efficient and feasible, we select an RKHS space $\mathcal{H}$ defined on a compact metric space $\Omega$ with reproducing kernel $k(\cdot,\cdot)$. According to the Theorem 3 of [11], the properties of the Lemma 3.1 are held when the kernel of $\mathcal{H}$ is universal. Let $\mathcal{F}$ be a unit ball in RKHS $\mathcal{H}$. Then $\operatorname{MMD}[p,q;\mathcal{F}]=0$ if and only if $p=q$. The target of the optimization problem (3.10) can be rewritten as
$\underset{f\in\mathcal{F}}{\sup}\left(\mathbb{E}_{\bm{x}\sim\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k}}}[f(\bm{x})]-\mathbb{E}_{\bm{y}\sim\pi^{\mathcal{S}}_{\bm{X}_{k}\mid\bm{Y}_{1:k}}}[f(\bm{y})]\right).\\$ (3.11)
The formula (3.11) is also called Maximum Mean Discrepancy. The MMD is a reasonable measure of the difference between two distributions. When MMD equals zero, two distributions are identical. With the reproducing property of RKHS, we do not need to compute the expectation of a specific function $f$ directly. It suffices to compute (3.11) by (3.12) in the following lemma.
@@
(Lemma 5 in [11]) Given $\bm{x}$ and $\bm{x}^{\prime}$ independent variables with distribution $p$ , $\bm{y}$ and $\bm{y}^{\prime}$ independent variables with distribution $q$. Assuming that the expectation of feature mapping $\phi(\cdot)$ of RKHS $\mathcal{H}$ exists. Then
$\operatorname{MMD}\limits^{2}[p,q;\mathcal{F}]=\mathbb{E}_{\bm{x},\bm{x}^{\prime}\sim p}[k(\bm{x},\bm{x}^{\prime})]-2\mathbb{E}_{\bm{x}\sim p,\bm{y}\sim q}[k(\bm{x},\bm{y})]+\mathbb{E}_{\bm{y},\bm{y}^{\prime}\sim p}[k(\bm{y},\bm{y}^{\prime})].$ (3.12)